Find the solution set of x1+2x2-3x3+x4=0

3X1-X2+5X3-X4=0

2X1+X2+X4=0


1
Expert's answer
2022-05-30T16:28:36-0400

Write down the augmented matrix


"A=\begin{pmatrix}
 1 & 2 & -3 & 1 &| & 0 \\
 3 & -1 & 5 & -1 &| & 0 \\
 2 & 1 & 0 & 1 &| & 0 \\
\end{pmatrix}"

"R_2=R_2-3R_1"


"\begin{pmatrix}
 1 & 2 & -3 & 1 &| & 0 \\
 0 & -7 & 14 & -4 &| & 0 \\
 2 & 1 & 0 & 1 &| & 0 \\
\end{pmatrix}"

"R_3=R_3-2R_1"


"\begin{pmatrix}
 1 & 2 & -3 & 1 &| & 0 \\
 0 & -7 & 14 & -4 &| & 0 \\
 0 & -3 & 6 & -1 &| & 0 \\
\end{pmatrix}"

"R_2=-R_2\/7"


"\begin{pmatrix}
 1 & 2 & -3 & 1 &| & 0 \\
 0 & 1 & -2 & 4/7 &| & 0 \\
 0 & -3 & 6 & -1 &| & 0 \\
\end{pmatrix}"

"R_1=R_1-2R_2"


"\begin{pmatrix}
 1 & 0 & 1 & -1/7 &| & 0 \\
 0 & 1 & -2 & 4/7 &| & 0 \\
 0 & -3 & 6 & -1 &| & 0 \\
\end{pmatrix}"

"R_3=R_3+3R_2"


"\begin{pmatrix}
 1 & 0 & 1 & -1/7 &| & 0 \\
 0 & 1 & -2 & 4/7 &| & 0 \\
 0 & 0 & 0 & 5/7 &| & 0 \\
\end{pmatrix}"

"R_3=7R_3\/5"


"\begin{pmatrix}
 1 & 0 & 1 & -1/7 &| & 0 \\
 0 & 1 & -2 & 4/7 &| & 0 \\
 0 & 0 & 0 & 1 &| & 0 \\
\end{pmatrix}"


"R_1=R_1+R_3\/7"


"\begin{pmatrix}
 1 & 0 & 1 & 0 &| & 0 \\
 0 & 1 & -2 & 4/7 &| & 0 \\
 0 & 0 & 0 & 1 &| & 0 \\
\end{pmatrix}"

"R_2=R_2-4R_3\/7"


"\begin{pmatrix}
 1 & 0 & 1 & 0 &| & 0 \\
 0 & 1 & -2 & 0 &| & 0 \\
 0 & 0 & 0 & 1 &| & 0 \\
\end{pmatrix}"

Take "x_3=t, t\in \R."

Then "x_1=-t, x_2=2t, x_3=t, x_4=0, t\in \R"



"\{-t, 2t, t, 0\}, t\in \R"